OpenStudy (graciek):

what do you do to check wheater a number is rational or irrational? in your explanation use an example of an irrational and a rational number.

OpenStudy (anonymous):

a number which can be written in the form p/q where p and q are integers. e.g \[2=\frac{ 2 }{ 1 },-3=\frac{ -3 }{ 1 },1.3=\frac{ 13 }{ 10 },0.666.....=\frac{ 2 }{ 3 }\] are rational numbers. \[\sqrt{7},\sqrt{2},\pi ~etc.~are~irrational~numbers.\]

Miracrown (miracrown):

Well by definition, rational numbers are fractions. that is, numbers that can be expressed as a ratio of two integers \[\frac{ 5 }{ 4 }\] or \[\frac{ 8 }{ 3 }\] or \[\frac{ 239487293 }{ 238492837982 }\] these are all rational numbers Usually, irrational numbers are numbers that involve square roots, like: \[\sqrt2\] or \[\sqrt20\] irreducible square roots are irrational numbers Other examples of irrational numbers are the famous constants, pi and e these are also irrational - cannot be expressed as a ratio of two integers

OpenStudy (graciek):

thanks

Miracrown (miracrown):

yw :)